I am confused on this question:
complete the ordered pairs so they are solutions of the equation 6x-5y=-120?
The answers are (_,0) then (0,_) and then (_,6).

Answer by addingup(3677)

6x-5y=-120
Solve for:
(_,0), (0,_), (_,6)
The first value is always x and the second value always y.
Let's find the first x in order to complete the first pair:
6x-5y = -120, and y = 0:
6x-5(0) = -120 and -5*0 = 0 so now we have:
6x = -120
x = -60
Now we know the first pair is (-60, 0)
``````````````````````````````````````````````````````````````````
Solve for y on the second pair:
6(0)-5y = -120
-5y = -120
y = -24
This gives us our second pair: (0,-24)
``````````````````````````````````````````````````````````````````
And on to the third:
6x-5y = -120
6x-5(6) = -120
6x-30 = -120
6x = -90
x = -15
So now we have our last pair:
(-15, 6)